# New Online Communities: Graph Deep Learning on Anonymous Voting Networks to Identify Sybils in Polycentric Governance

### v4

@@ 6.3.3 Graph Convolutional Neural Network (GCNN)
With the voting graph, we engineer learnable features and train a Graph Convolutional Neural Network (GCNN) to create vector embeddings for each node. Since there are any number of possible custom GCNNs tailored for node classification tasks on graphs, with a multitude of interchangeable layers for various tasks, we designed the mechanism inductively with iterative testing against the dataset.
@@
For our method, the first few steps are relatively straightforward. First, construct a voting graph where known (labelled) voters and re-used wallets are combined into single nodes with multiple edges to proposal nodes (we use a multi- or hyper-graph to accommodate multiple edges from each node). Second, we engineer features to learn higher dimensional vector embeddings on all nodes. We use a simple reconstruction loss (Mean Squared Error or MSE) to capture node features for training the GCNN (we cannot use a more conventional classification training loss like cross-entropy because we do not know the sybil classes a priori; instead, we use the GCN only to create the embeddings). Third, we use the node embeddings to conduct a k-means vector clustering on unknown nodes. We then propagate the most common label associated with each cluster to relabel the actual, lower dimensional nodes to form the sybil cluster. Finally, we take the relabelled similarity graph and combine nodes with their labels in each cluster to form a higher dimensional, reduced-size graph. The resulting graph is an inductive prediction of sybil nodes.
@@
The deep learning method was developed with the Python libraries NetworkX and RAPIDS AI to construct the graphs, the PyTorch Geometric library for the GCNN, and the Facebook AI Similarity Search (FAISS) for the fast embeddings clustering. Whenever possible, GPU accelerated libraries were used, although speed and memory tradeoffs quickly become a practical data analysis challenge when dealing with large graphs requiring several hundred gigabytes of memory for large matrix multiplications during learning. As an inductive method, decisions to use a particular library, feature, or machine learning algorithm were made pragmatically and tested against the dataset.
@@
We develop features and construct a typical train/test/validation split for the forward pass of our GCNN. The first layer of the forward pass is fully connected, capturing node features and incorporating voting edges with their ‘power’ attributes. The second layer integrates edge and node features with a sequential multi-layer perceptron (MLP) using ReLu activation, which is then fed into a convolutional layer with mean feature aggregation. The third layer processes temporal sequences engineered for a Long Short-Term Memory (LSTM) layer. Finally, the fourth layer uses a multi-headed Graph Attention Network (GAT) to better learn and search edge connectivity. Each of these layers has tuneable hyperparameters that were set by conducting a grid search. After each forward pass and backpropagation to set training weights, training loss (MSE) was calculated, visualized, and assessed. The resulting tensors are normalized on zero and were checked for sufficient numeric variability.
The four layer GCNN to identify sybils:
@@ 6.3.4 Fast Vector Embeddings Search using FAISS
Once we have learned higher dimensional voter and voting representations for all (known and unknown) nodes, we can cluster similar vector embeddings, which we initially learned using the GCNN. In deciding on our method, we also considered other similarity metrics like Euclidian distance, cosine similarity, and Jaccard similarity, but computing many-to-many similarity searches on large matrixes is very slow $O(n^{2}\cdot d)$. Thankfully, the Facebook AI Similarity Search (FAISS) Python library provides extremely fast and easily scaled search and clustering options (including GPU processing). FAISS offers several search index options for improving efficiency, including approximate search methods that can scale to billions of elements. For our purposes, the L2 index was quick enough and provides exact results (for much larger datasets, FAISS provides approximate methods that are substantially quicker). We use the included k-means clustering method to find sybil clusters, which uses centroids in the high dimensional embeddings to cluster and return neighbours. We normalize the clusters by dropping any remaining singletons and large clusters (mean + 1 standard deviation).
@@
We then inspect each cluster and propagate the most common label across the cluster to construct a “similarity” graph. Finally, we construct a reduced “clustered” graph by combing clustered nodes (and reconstructing edges).
@@ 7   Results
We have conducted numerous graph analysis experiments on DAO governance voting with a range of graph sizes, learnable features, specialized convolutional layers, and training and optimization algorithms. We also tested different clustering approaches including DBMScan, spectral clustering, and unsupervised methods; we deployed supervised classification methods with the labelled subgraph; and we considered the problem from different graph analysis viewpoints (such as considering sybil identification as a problem of link prediction using graph autoencoders, and as a community detection problem using Girvan-Newman and Louvain). While many of these methods initially looked promising, they ultimately failed for a variety of reasons.
The most common reason for experimental failure was the lack of edge connectivity on sybils, which is precisely what provides anonymity. However, our task of identifying sybils is easier than complete deanonymization and by approaching the problem as a k-anonymous graph signal we might achieve what we want anyway, which is the ability to defend against attackers at scale but not reveal individual identities. This is an easier, but not easy problem. Since we should expect sybils to be singletons, also like voting spam, we cannot effectively cluster their lower dimensional structure, nor can we use a labelled subgraph of known users for supervised learning, and nor will we be able to detect communities or cliques. Instead, we defined sybils as clusters of similar higher dimensional vector node embeddings. And since we form clusters using a k-means clustering method on higher dimensional embeddings, singletons pose no particular challenge for our method.
Tested against the snapshot.org dataset, our method typically reduces the original voting graph by approximately 2-5% (subject to hyperparameters). The method produces a voting graph with real IDs that can be inspected for further forensic analysis or visualized (see Figures 13 and 14). Moreover, we can produce a few key sociometric measurements to begin the process of constructing community baselines for further research (see Table 2).
@@ 8   Future Research
To improve and further validate our inductive method of sybil identification we need benchmarks from other online communities. Moreover, more sophisticated and further development of machine learning features would enhance the method. Finally, despite the computational demands required to ‘follow the money,’ using this direct method to identify sybil chains is the most reliable signal we have, and if conducted at scale the method would provide an invaluable graph signal for supervised learning.
In addition to refining current methods, it is worth exploring other deep learning techniques. To adapt the method for detecting voting spam, we could instead detect anomalies within the node embeddings. Thus, instead of identifying similar nodes, this approach would search for outliers to indicate spam activity. In coin voting contexts, where the integrity of each vote is tied to a digital token, distinguishing between voting anomalies (spam) and similarities (sybils) is critical for safeguarding the democratic processes within DAOs.
Finally, deep learning methods can be exploited by malicious users. Making deep learning methods resilient to attack is an active area of research and a practical requirement for DAOs, which must protect their digital commons from theft and inappropriate extraction. Examples of attacks against GCNNs include adversarial attacks that manipulate input data to deceive models, model inversion attacks that reveal sensitive information, and poisoning attacks that corrupt training data. Developing robust models that can detect and resist these attacks is crucial. Additionally, understanding and mitigating the ethical implications of AI and ML in DAO governance, such as bias in decision-making processes, must be a priority. These challenges open up diverse research avenues, ranging from developing new algorithms for anomaly detection to ethical frameworks for AI deployment in decentralized governance structures.

### v6

@@ 5.3.4 Graph Convolutional Neural Network (GCNN)
With a large voting graph, we engineer learnable features and train a Graph Convolutional Neural Network (GCNN) to create vector embeddings for each node and its edge connections. Since there are any number of possible custom GCNNs tailored for self-supervised node classification tasks on graphs, each with a multitude of interchangeable layers for various tasks, we designed the mechanism inductively with iterative testing against the dataset. We proceed experimentally and use any tools and techniques that work well. For instance, above we used a deep feedforward neural network architecture to learn the parameters and then clustered with a Random Forest, which also worked well. In general, as Simon Prince notes [47], most overparameterized deep models work well, despite vast differences in algorithm architecture and mathematical approach.
@@
For our method, the first few steps are relatively straightforward. First, construct a voting graph where known (labelled) voters and re-used wallets are combined into single nodes with multiple edges to proposal nodes (we use a multi- or hyper-graph to accommodate multiple edges from each node). Second, we engineer features to learn higher dimensional vector embeddings on all nodes (this is the encoder). We use a simple reconstruction loss (Mean Squared Error or MSE) to capture node features for training the GCNN (we do not use a more conventional classification training loss like cross-entropy because we do not know the Sybil classes a priori). Third, we use the node embeddings to conduct a k-means vector clustering on unknown nodes (this is the decoder). We then propagate the most common label associated with each cluster to relabel the actual, lower dimensional nodes to form the Sybil cluster (this is also a type of MaxPooling layer for dimensionality reduction). Finally, we take the relabelled similarity graph and combine nodes with their labels in each cluster to form a higher dimensional, but reduced-size graph (upscaling). The resulting graph is an inductive prediction of Sybil nodes.
@@
The deep learning method was developed with the Python libraries NetworkX and RAPIDS AI to construct the graphs, the PyTorch Geometric library for the GCNN, and the Facebook AI Similarity Search (FAISS) for the fast embeddings clustering. Whenever possible, GPU accelerated libraries were used, although speed and memory tradeoffs quickly become a practical data analysis challenge when dealing with large graphs requiring several hundred gigabytes of memory for large matrix multiplications during learning. This is known as the graph expansion problem and can be controlled by neighborhood sampling or graph partitioning (batch normalization). As an inductive method, decisions to use a particular library, feature, or machine learning algorithm were made pragmatically and tested against the dataset, but as discussed below, the GCNN encoder-decoder approach offers some distinct conceptual advantages over other approaches.
@@
We develop features and construct a typical train/test/validation split for the forward pass of our GCNN. The first layer of the forward pass is fully connected, capturing node features and incorporating voting edges with their ‘power’ attributes. The second layer integrates edge and node features with a sequential multi-layer perceptron (MLP) using ReLu activation, which is then fed into a convolutional layer with mean feature aggregation. The third layer processes temporal sequences engineered for a Long Short-Term Memory (LSTM) layer. Finally, the fourth layer uses a multi-headed Graph Attention Network (GAT) to better learn and search edge connectivity. Due to the final attention layer, this encoder can be considered a transformer with residual connections. (Due to dataset availability limitations, the wallet fingerprint signal was not used.)
Each of these layers has tuneable hyperparameters that were set by conducting a grid search. After each forward pass and backpropagation to set training weights, training loss (MSE) was calculated, visualized, and assessed. The resulting tensors were normalized on zero and were checked for sufficient numeric variability.
Summarizing, the four layer GCNN to identify Sybils:
@@ 5.3.5 Fast Vector Embeddings Search using FAISS
Once we have learned higher dimensional voter and voting representations for all (known and unknown) nodes, we can cluster similar vector embeddings (an encoding task), which we initially learned using the GCNN. In deciding on our method, we also considered other similarity metrics like Euclidian distance, cosine similarity, and Jaccard similarity, but computing many-to-many similarity searches on large matrixes is very slow $O(n^{2}\cdot d)$. Thankfully, the Facebook AI Similarity Search (FAISS) Python library provides extremely fast and easily scaled search and clustering options (including GPU processing) for vector embeddings. FAISS offers several search index options for improving efficiency, including approximate search methods that can scale to billions of elements. For our purposes, the L2 index was quick enough and provides exact results (for much larger datasets, FAISS provides approximate methods that are substantially quicker). We use the included k-means clustering method to find Sybil clusters, which uses centroids in the high dimensional embeddings to cluster and return neighbours. We normalize the clusters by dropping any remaining singletons and large clusters (mean + 1 standard deviation).
@@
We then inspect each cluster and propagate the most common label across the cluster to construct a “similarity” graph (downsampling). Finally, we construct a reduced “clustered” graph by combing clustered nodes (and reconstructing edges) (upsampling).
@@ 6   Results
We have conducted numerous graph analysis experiments on DAO governance voting with a range of graph sizes, learnable features, specialized convolutional layers, and training and optimization algorithms. We also tested different clustering approaches including DBMSCAN and other unsupervised methods; we deployed supervised classification methods with the labelled subgraph of true (Ethereum Name Service) identities; and we considered the problem from different graph analysis viewpoints (such as considering Sybil identification as a problem of link prediction or as a community detection problem). While many of these methods initially looked promising, many failed for a variety of reasons; most commonly, approaches cannot rely on (k$>$1) edge connectivity, which is precisely what provides anonymity.
However, our task of identifying Sybils is easier than complete deanonymization and by approaching the problem as a k-anonymous graph signal we might achieve what we want anyway, which is the ability to defend against attackers at scale but not reveal individual identities. This is an easier, but not easy problem (avoiding Douceur’s [14] strict definition of a Sybil). Since we should expect Sybils to be singletons, also like voting spam, we cannot effectively cluster their lower dimensional structure, nor can we use a labelled subgraph of known users for supervised learning, and nor will we be able to detect communities or cliques. Instead, we defined Sybils as clusters of similar higher dimensional vector node embeddings. And since we form clusters using a k-means clustering method on higher dimensional embeddings, singletons pose no particular challenge for our method.
Graph convolutional neural networks offer several advantages for identifying Sybils. First, network topology is informative but difficult to learn using traditional methods. This is especially true for very large graphs. However, since convolutional networks aggregate information from nearby nodes, and therefore inductively bias these relationships, learning is simplified (the graph signal is downsampled), and since parameters are shared across convolutional layers (lowering memory requirements), large graphs are less difficult to learn. Moreover, our encoder-decoder approach takes cues from Natural Language Processing (NLP), which must accommodate a large number of token inputs (precluding deep, fully connected networks). Our use of a multi-head graph attention layer produces node embeddings with n-deep neighborhood information (this is also known as a transformer), thereby increasing network depth but with tractable sparsity. Similarly, the Long-Short Term Memory (LSTM) algorithm used to learn time series data can be considered an earlier version of the transformer model. More broadly, transformer models, although still rarely applied to graph problems, offer significant untapped potential for network deanonymization and have already proven useful for image processing and other discrete structural problems.
Tested against the snapshot.org dataset, our method typically reduces the original voting graph by approximately 2-5% (subject to hyperparameters). Quite obviously, performance could be improved and validation measurements need to be established. The method produces a voting graph with real IDs that can be inspected for further forensic analysis or visualized (see Figures 24 and 25). Moreover, we can produce a few key sociometric measurements to begin the process of constructing community baselines for further research (see Table 3).
@@
Finally, we used the trained embeddings to generatively produce voting samples, which we stimulated with varying Sybil signals. This voting simulation accurately reflects the temporal dynamics and evolution of online communities by generating samples from high dimensional embeddings representing specific voting behaviours. The introduction of Sybil signals reveals potential sources of resource congestion and over extraction.
DRAFT: INSERT DIAGRAMS
@@ 7   Future Research
To improve and further validate our inductive method of Sybil identification we need benchmarks from other online communities. Moreover, more sophisticated and further development of machine learning features would enhance the method. Finally, despite the challenges required to ‘follow the money,’ using this direct method to identify Sybil chains is the most reliable signal we have, and if conducted at scale the method would provide an invaluable graph signal for supervised learning.
In addition to refining current methods, it is worth exploring other deep learning techniques. To adapt the method for detecting voting spam, we could instead detect anomalies within the node embeddings. Instead of identifying similar nodes, searching for outliers might indicate spam activity (perhaps this might be exploited using transfer learning). Or, since the GCNN trains an embeddings model, generative tasks like link prediction or economic simulation might be possible.
Finally, deep learning methods can be exploited by malicious users. Making deep learning methods resilient to attack is an active area of research and a practical requirement for DAOs, which must protect their digital commons from theft and inappropriate extraction. Examples of attacks against ML models include adversarial attacks that manipulate input data to deceive models, model inversion attacks that reveal sensitive information, and poisoning attacks that corrupt training data. Developing robust models that can detect and resist these attacks is crucial and should be part of any production environment. Additionally, understanding and mitigating the ethical implications of AI and ML in DAO governance, such as bias in decision-making processes, must be a priority. These challenges open up diverse research avenues, ranging from developing new algorithms for anomaly detection to ethical frameworks for AI deployment in decentralized governance structures.